# Compositional disorder in a multicomponent non-reciprocal mixture: stability and patterns

### v2

@@ Abstract
The mean compositions of individual components can be tuned to control phase behavior in number-conserving passive mixtures. In this work, we investigate the role of variable average density in a system of infinitely many non-reciprocally interacting scalar densities, within the framework of the multi-species non-reciprocal Cahn-Hilliard (NRCH) model. Rather than focusing on specific parameter choices, we study ensembles of systems where the inter-species interaction coefficients and average densities are sampled from probability distributions. We show that non-reciprocity stabilizes the homogeneous mixed state even in the presence of compositional disorder, when compared to the corresponding reciprocal system. Using random matrix theory, we derive a general condition for the onset of spinodal instability, which we verify through simulations. Finally, we illustrate the connection between the statistics of the most unstable eigenvalue and the emergent nonlinear dynamics.
@@ I Introduction
An important question that has driven intense research is – how does a cell achieve self-organization and regulation? A widely accepted route, applicable in many scenarios, is that components in the cell cytoplasm organize into compartments enriched in a specific group of chemicals simply through thermodynamic phase separation [1]. In the past decade, liquid-liquid phase separation has been implicated in the formation of membrane-less organelles [2, 3, 4, 5], regulation of stable structures such as the nucleolus [6, 7, 8], and adaptation to adverse external stimuli in stress granules [9, 10]. It is striking that this intuitively appealing idea surfaced for the first time at the end of the nineteenth century [11, 12] and again in the context of proto-cells [13]. The identification of phase separation as the dominant mechanism opens up new possibilities for regulating condensate formation through pathways accessible in passive systems. These include controlling pH [14], modulating chemical activity [15], coupling to an active fluid [16], and, most notably, altering the mean density of the constituents [17]. The latter presents a particularly promising strategy for regulating phase separation during droplet ripening in the early development of the C. elegans embryo [18].
@@
Figure 1: Role of compositional disorder in multi-species NRCH (a) The average composition of each species is fixed to a randomly chosen value. The dynamical matrix determining the linear stability of the homogeneous mixed state is the sum of a random Gaussian matrix (top) and a diagonal matrix whose entries are determined by the mean density of each species (bottom). In (a) the diagonal entries are drawn from uniform distribution and each color shows a different random value. (b) Three other distributions are chosen to illustrate our main results - the associated probability distribution functions and their name is mentioned. (c) To illustrate the effect of compositional disorder on the dynamics, we run simulations in one dimensional space keeping $\mathbb{M}$ the same in all three cases. The top and the bottom evolve to a state of spatiotemporal chaos in the steady state, while the middle evolves to a phase separated state.
@@
A description of processes within the cell cytoplasm has to incorporate two features: multiple participants and activity in a suitable form. It is well known that the cell interior is a heterogeneous multicomponent environment [19], a feature reflected in numerous studies on passive phase  [20, 21, 22, 23, 24]. However, the perspective that activity must be included in a system-specific form in the formation of bio-organelles is relatively recent [25]. In the field of active matter, active phase separation has been explored through several minimal models that typically involve either a single component—used to model motility-induced phase separation or coupling to chemical reactions [26, 27, 28, 29]—or mixtures with two or more non-reciprocally interacting components [30, 31].
With this background, we take a minimal approach where we model the cell as a multicomponent mixture with non-reciprocal inter-species interactions [32, 33, 34, 35, 36, 37, 38]. Non-reciprocity appears naturally in mixtures of active particles systems: whether polar [39] or scalar [30, 31], and also in quorum sensing mixtures [40, 41], or mixtures of Janus colloids [42]. The hallmark of non-reciprocal number conserving systems is that it combines features of pattern-forming systems [43, 44] and phase separating systems [45, 46]. Moreover, the dynamics can be tuned by the changing composition, thus transitioning from stable mixtures to traveling lattices through a conserved version of the Hopf-bifurcation [30].
We address a question that arises naturally - how does assigning a varying mean density to each species, hereafter referred to as compositional disorder, influence the stability and pattern formation within the system? In [36], it was shown that non-reciprocal activity in the form of asymmetric correlation between the interaction coefficients between the species stabilizes the homogeneous mixed state in comparison with the system with passive interactions. It means that a temperature like control parameter has to be lowered more to drive the active system to a patterned state.
@@
In this work we use random matrix theory [47, 48, 49, 50] to study the linear stability of the system as pioneered in the seminal works of Wigner [51], and May [52]. This approach has been successfully applied to a wide range of physical systems, including the stability of ecosystems with random interactions [53, 54, 55, 56, 57], and disordered systems such as spin glasses [58, 59, 60]. We adopt a statistical perspective, replacing the complexity of individual interaction networks with probabilistically sampled ensembles of systems [21, 61, 62]. Within this framework, the mean composition is treated as a stochastic variable drawn from a distribution, as recently investigated in passive systems [63]. The linearized dynamics of the multi-species NRCH model reduces to studying the stability of the asymmetric matrices with diagonal disorder as represented in Fig. 1(a) facilitating the statistical estimation of the onset of the spinodal instability [64], and the possible phases in this system.
@@
We show that for any choice of compositional disorder, the non-reciprocally interacting system is more stable than the passive one. The theoretical predictions are supported by the numerical solution of the nonlinear model for some choices of the distribution for the average densities shown in Fig. 1(b). We demonstrate the role of complex eigenvalues in producing diverse dynamics that are not permitted in the equilibrium counterpart. The paper is organized as follows: we introduce the theoretical framework of the multi-species NRCH model in Section II. In Section III, we discuss some known results to determine the linear stability of a system with compositional disorder in which case it is essential to consider random Gaussian matrices whose diagonal matrices have been drawn from a different distribution. In the next section, we present some general properties of the dynamical matrix leading to the results in section III. We discuss the varied dynamics seen in the model in section V and end with a summary of the main results in the paper in VI.
@@
We adopt the framework presented in [36] for a multicomponent active mixture and consider $N$ individually conserved active species. These species interact via pairwise non-reciprocal interactions originating from the active nature of the system, and reciprocal interactions driven by a free energy. The scalar field $\phi_{a}({\bm{r}},t)$ represents the number density of the $a$-th species. The number conserving densities obey gradient dynamics of the form $\partial_{t}\phi_{a}=\Gamma_{a}\nabla^{2}\mu_{a}$ with mobility $\Gamma_{a}$ and active chemical potentials $\mu_{a}$. $\mu_{a}$ consists of two parts: the first is derivable from a free energy, and the second represents the non-equilibrium active contribution. The passive part is derived from the free energy functional $F=\int\mbox{d}^{d}\bm{r}f$, where $f$ is the free energy density which promotes macroscopic phase separation. We use a minimal polynomial form for $f$ as shown below
$f$ $=$ $\sum_{a}\left[\frac{1}{2}\Theta_{a}\phi_{a}^{2}+\frac{s_{a}}{4}\phi_{a}^{3}-\frac{K}{2}|\bm{\nabla}\phi_{a}|^{2}\right]$ (1)
@@
$\chi_{ab}$ is Flory-Huggins type interaction parameter between two species $a$ and $b$. To make progress, we have made some assumptions to make the problem more tractable while still retaining enough complexity. The interfacial tension is assumed to be identical for all species and equal to $K$ to focus on effects that are distinct from a Turing type instability that appears when the tensions are chosen to be different [65]. The coefficient of self-interaction $\Theta_{a}=(1+\Theta)$ is the same for all species, akin to an external tuning parameter like the temperature of the system that controls the transition to a patterned state. We now rescale time, space, and the scalar fields as $t\to{K}/{\Gamma}t$, $x\to\sqrt{K}x$, and $\phi_{a}\to\phi_{a}/\sqrt{s_{a}}$, to recast the equations as follows
@@
$-\nabla^{4}\phi_{a},\,\,\int\mbox{d}^{d}\bm{r}\phi_{a}(\bm{r},t)=\phi_{a0}.$ (2)
We consider a matrix $\mathbb{M}$ of general interactions whose coefficients are drawn from a Gaussian distribution of zero mean and variance $\sigma$.
$M_{ab}(\chi,\alpha)=\frac{\bar{\alpha}}{\sqrt{2N(\bar{\alpha}^{2}+\bar{\chi}^{2})}}A_{ab}+\frac{\bar{\chi}}{\sqrt{2N(\bar{\alpha}^{2}+\bar{\chi}^{2})}}S_{ab},$ (3)
where $\mathbb{A}$ and $\mathbb{S}$ are fully asymmetric and symmetric matrices respectively, i.e. $A_{ab}=-A_{ba}$, and $S_{ab}=S_{ba}$. Eq. (3) is written such that the variance of the coefficients is $\sigma^{2}$ and the correlation between the diagonal partners determines the level of non-reciprocity in the system, which we call $\alpha$. The properties of $\mathbb{M}$ are summarised as follows
$\langle M_{ab}M_{ab}\rangle=\frac{\sigma^{2}}{2N},\,\,M_{aa}=0,$
$\langle M_{ab}M_{ba}\rangle=\frac{\sigma^{2}}{2N}(1-2\alpha^{2}),\,\,\alpha=\frac{\bar{\alpha}}{\sqrt{\bar{\alpha}^{2}+1}}.$ (4)
@@
To determine the linear stability of the system we substitute $\phi_{a}=\phi_{a0}+\delta\phi_{a}$ in Eqs. (II) and to obtain the linearised dynamics around the homogeneous state
$\partial_{t}\delta\phi_{a}=-q^{2}\sum_{b}\left[M_{ab}+(1+\Theta+3\phi^{2}_{a0})\delta_{ab}\right]\delta\phi_{b}.$ (5)
Defining a diagonal matrix $C_{aa}=3\phi^{2}_{a0}$, whose off-diagonal entries are zero, it is clear from Eq. (5) that the linear stability of the system is determined by the eigenvalues of the matrix $\mathbb{D}=\mathbb{M}+\mathbb{C}$. The effect of varied composition thus enters the dynamical matrix as diagonal disorder through the diagonal matrix $\mathbb{C}$ whose diagonal elements $\bar{\phi}$ are drawn from a distribution $P(\bar{\phi})$. The boundary of the spinodal instability is defined as a part of the multi-dimensional parameter space where at least one of the N eigenvalues is unstable. The onset of the spinodal instability is thus determined by the eigenvalue $\lambda_{0}$ of $\mathbb{D}$ with the smallest real part. With this motivation, we will first discuss the onset of pattern formation in the system.
Figure 2: Spinodal instability in theory and simulations: (a) The boundary of linear stability is plotted with a solid line in the $\Theta-\alpha$ plane, for a fixed set of parameters for the discrete bimodal distribution, see Eq. (13). The predictions are tested using simulations with $50$ realisations of $\mathbb{M}$. The colour of the markers reflects the ensemble and space-averaged order parameter $\Delta$ [see Eq. (9)], which vanishes outside the spinodal shown by a light hue and changes discontinuously to a finite value inside the unstable region, as marked by a darker hue. The grey line is the same for identical mean composition of the densities. The plots show that the relative stability depends both on the distribution and on $\alpha$ (b) Same as panel (a) for the exponential distribution.
@@ III Spinodal
The most unstable eigenvalue $\lambda_{0}$ determines the threshold value of self-interaction or temperature $\Theta_{\rm sp}$ of the linearly stable system,
@@
where $\{D\}$ denotes the set of parameters that characterize the distribution $P$. The statistics of $\lambda_{0}$ are determined by the properties of the dynamical matrix $\mathbb{D}$. For $\Theta$ lower than $\Theta_{\rm sp}$, the homogeneous state is unstable and patterns are formed in the steady state.
For $P(\bar{\phi})=\delta(\bar{\phi}-3\phi_{0}^{2})$, i.e. if the mean density of each species is identical, $\lambda_{0}$ is given by the eigenvalues of $\mathbb{M}$ shifted by a constant factor [36].
@@
Eq. (7) implies that for a fixed $\phi_{0}$, $\Theta_{\rm sp}$ is smaller for the active system in comparison to the passive system ($\alpha=0$). The control parameter $\Theta+1$ is synonymous with the strength of self-interaction. If all entries of the matrix $\mathbb{M}$ were set to zero, the non-interacting system would not phase separate for $\Theta>-1$. $\Theta<-1$ implies attractive interactions between particles of the similar type leading to phase separation. Eq. (6) shows that random passive interactions generate stronger effective attractive interactions between the species, such that to tune the system to a homogeneous state, it is necessary to increase intra-species repulsive interactions. For a fully non-reciprocal system, $\alpha=1$, $\Theta_{\rm sp}$ decreases to $-1$ which is the threshold for the non-interacting system. The non-reciprocal counterpart is thus more stable than the reciprocal system. The factor of unity in Eq. (6) simply sets $\Theta_{\rm sp}$ for passive interactions to zero which serves as a good reference point for comparisons.
@@
The slope $m$ and $l_{0}$ depend only on the parameters $\{D\}$ characterizing the distribution $P$. $l_{0}$ determines the stability of the passive system, $\alpha=0$, in the presence of compositional disorder. To compare the stability of the state with compositional disorder, with $l_{0}>(-3\phi_{0}^{2}+1)$, compositional disorder stabilizes the homogeneously mixed state. The sign of the ‘slope’ $m$ between $\Theta_{\rm sp}$ and $\alpha^{2}$ determines the stability of the active system. For $m<0$ the active system is more stable than the passive one.
In Fig. 2 we have plotted the theoretical prediction for $\Theta_{\rm sp}$ for the bimodal and exponential distributions described in IV, see Eqs. (13) and (14). To test the predictions of linear stability analysis, we solve Eq. (II) numerically. We start our simulations with uniform composition $\phi_{a0}$ such that $3\phi_{a0}^{2}$ is drawn from either the bimodal or the exponential distributions. The simulations are initiated with random Gaussian perturbations. For other details on system size, methods used, see the Appendix IV. The results of the simulation are shown in Fig. 2, where the markers denote the points where the simulations are carried out and the colour of the markers carries information about the magnitude of the order parameter.

### v3

# Compositional disorder in a multicomponent nonreciprocal mixture: stability and patterns

@@ Abstract
In passive phase-separating mixtures, the average density of each component can be tuned to control the composition of the coexisting bulk phases. This concept extends to active systems. For example, in a nonreciprocal mixture of two species, changing the average density of either species alters the qualitative nature of the travelling patterns that emerge in the steady state. In this paper, we build on the existing framework of the multi-species nonreciprocal Cahn–Hilliard (NRCH) model to examine the influence of compositional disorder in a multicomponent active system. Specifically, we allow each scalar field in the mixture to have a distinct average density and analyze the implications of this generalized setting. Focusing on ensembles of systems in which the inter-species interaction coefficients and the average densities of each species are both sampled from probability distributions, we show that nonreciprocity stabilizes the uniformly mixed state, even in the presence of compositional disorder, when compared to the corresponding reciprocal system. Using random matrix theory, a general condition for the onset of the spinodal instability is derived and verified through simulations. Finally, the connection between the statistics of the most unstable eigenvalue and the emergent nonlinear dynamics is illustrated.
@@ I Introduction
An important question that has driven intense research is – how does a cell achieve self-organization and regulation? A widely accepted route, applicable in many scenarios, is that the constituents of the cell cytoplasm organize into compartments enriched in a specific group of chemicals simply through thermodynamic phase separation [1, 2]. In the past decade, liquid-liquid phase separation has been implicated in the formation of membrane-less organelles [3, 4, 5, 6], regulation of stable structures such as the nucleolus [7, 8, 9], and in strategies to adapt to adverse external stimuli in stress granules [10, 11]. It is striking that this intuitively appealing idea that has recently received widespread acceptance, surfaced for the first time at the end of the nineteenth century to explain the shape of sea-urchin protoplasm [12, 13] and again in the context of proto-cells which might have paved the way for the origin of life [14]. The identification of phase separation as the dominant mechanism opens up new possibilities for regulating condensate formation through pathways accessible in passive systems. These include controlling pH [15], modulating chemical activity [16], applying a chemical gradient [17], coupling to an active fluid [18], and, most relevant to this work, altering the relative mean density of the constituents [19]. The exploration of the last pathway, namely, controlling the total availability of a particular chemical species in an active phase-separating mixture, is the main topic of this paper.
@@
Figure 1: Role of compositional disorder in multi-species NRCH: (a) The average composition of each species is fixed to a value that is drawn from a distribution. The dynamical matrix determining the linear stability of the homogeneous mixed state is the sum of a random Gaussian matrix (top) and a diagonal matrix whose entries are determined by the mean density of each species (bottom). In (a) the diagonal entries are drawn from a uniform distribution and each color shows a different random value. (b) Three other distributions are chosen to illustrate our main results - the associated probability distribution functions and their name is mentioned. (c) To illustrate the effect of compositional disorder on the dynamics, we run simulations in one dimensional space keeping $\mathbb{M}$ the same in all three cases. The top and the bottom evolve to a state of spatiotemporal chaos in the steady state, while the middle evolves to a phase separated state.
@@
A description of processes within the cell cytoplasm has to incorporate two features: multiple participants [20] and activity in a suitable form. A recent protein database [21] lists more than two thousand relevant types of proteins in membrane-less organelles that are collected from hundreds of papers. The heterogeneity at the cellular level is reflected in numerous studies on passive phase separation [22, 23, 24, 25, 26]. However, the perspective that active processes must be included in a suitable, system-specific form in a theoretical description of bio-condensates is relatively recent [27, 28], such that the properties of active mixtures constitute a new area of research. In the field of active matter, active phase separation in a single component has been explored in several minimal models, such as motility-induced phase separation [29, 30], chemically active mixtures [31, 32], and binary mixtures of nonreciprocally interacting components [33, 34].
With this background, we assume a reductionist approach where we model the cell as a multicomponent mixture with nonreciprocal inter-species interactions [35, 36, 37, 38, 39, 40, 41]. Nonreciprocity appears naturally in mixtures of active particles: whether the constituents have an orientation [42] or not [33, 34], in quorum sensing mixtures [43, 40], or mixtures of Janus colloids [44]. Nonreciprocity is a key ingredient that can be used to mimic behaviors typically associated with living, intelligent systems, such as predator–prey dynamics. Recent experimental findings have established the validity of some key theoretical predictions in polar mixtures [45], active solids [46, 47, 48], and a vibrated granular mixture of rod and spheres [49].
The hallmark of nonreciprocal number conserving systems is that it combines features of pattern-forming systems [50, 51] and phase separating systems [52, 53]. The latter aspect implies that the average density is an important physical quantity that can be tuned to alter dynamics, as shown explicitly for a binary nonreciprocal mixture where the system undergoes a transition from stable mixtures to traveling lattices via a conserved form of the Hopf bifurcation [33]. The role of composition is, however, largely unexplored in these systems, with a few exceptions in the binary case [33], and no studies in the multicomponent case.
In this work, we address a natural question: how does variability in the mean density of each species, hereafter referred to as compositional disorder, influence the stability and pattern formation in the system? Recent work [39] has shown that nonreciprocity, in the form of asymmetric correlations of inter-species interaction coefficients, stabilizes the homogeneous mixed state compared with a system that has only passive interactions. If the mean density of each species is identical, its effect can be subsumed into the definition of a temperature-like control parameter that can be tuned to drive pattern formation in the active system. The simplifying assumption, however, is invalid for biological condensates, as established in recent papers where it has been possible to measure the relative abundance of each component forming a condensate and show that it varies widely [54], implying that a realistic description of the underlying physics should account for this variability. In biological condensates the compositions are typically highly variable, for example in cancer cells [55] and in experiments using DNA chains to explore this diversity in a controlled manner [56].
@@
We use random matrix theory [57, 58, 59, 60] to study the linear stability of the system, as pioneered in the seminal works of Wigner [61] and May [62] in quantum mechanics and ecosystems, respectively. This approach has been successfully applied to a wide range of physical systems, including the stability of ecosystems with random interactions [63, 64, 65, 66, 67], and disordered systems such as spin glasses [68, 69, 70]. We adopt a statistical perspective, replacing the complexity of individual interaction networks with probabilistically sampled ensembles of systems [23, 71, 72]. Within this framework, the mean composition is treated as a stochastic variable drawn from a distribution, as recently investigated in passive systems [73]. The stability analysis of uniform mixtures of the multi-species NRCH model reduces to studying the eigenvalue spectrum of the asymmetric matrices with diagonal disorder (diagonal entries being pulled from a different, non-Gaussian distribution) as represented in Fig. 1(a). The approach facilitates the statistical estimation of the onset of the spinodal instability [74], and provides useful insights into the possible phases in this system.
@@
We show that for any choice of compositional disorder, the nonreciprocally interacting system is more stable than the passive one when one considers a system with infinitely many species with random interactions. The theoretical predictions are supported by the numerical solution of the full nonlinear model for a few chosen forms of the distribution for the average densities shown in Fig. 1(b). We demonstrate the role of complex eigenvalues, focusing on finite number of components, in producing diverse dynamics that are not permitted in the equilibrium counterpart. The paper is organized as follows: we introduce the theoretical framework of the multi-species NRCH model in Section II motivating why it is essential to consider random Gaussian matrices whose diagonal matrices have been drawn from a different distribution as illustrated in Fig. 1 (a). In Section III, we introduce a general condition pertaining to the linear stability of a system with compositional disorder. In Section IV, we present the analytical arguments leading to the results in section III. The varied dynamics possible in the model is explored in section V and we present our conclusions in VI.
@@
We follow the framework of [39] for multicomponent active mixtures and consider NN active species, each conserved individually. All species interact via pairwise nonreciprocal interactions originating from the active nature of the system [44, 38], and reciprocal interactions driven by a free energy. The scalar field $\phi_{a}({\bm{r}},t)$ represents the number density of the $a$-th species. The densities obey gradient dynamics of the form $\partial_{t}\phi_{a}=\Gamma_{a}\nabla^{2}\mu_{a}$ with mobility $\Gamma_{a}$ and active chemical potentials $\mu_{a}$. The chemical potentials $\mu_{a}=\mu_{a}^{(\rm eq)}+\mu_{a}^{(\rm ac)}$ have two parts each: the first is the functional derivative of a free energy $F$, and the second represents the non-equilibrium active contribution that is not related to a free energy. The free energy functional $F=\int\mbox{d}^{d}rf$, where $f$ is the free energy density which promotes macroscopic phase separation. We use a minimal polynomial form for $f$ as shown below
$f$ $=$ $\sum_{a}\left[\frac{1}{2}\Theta_{a}\phi_{a}^{2}+\frac{s_{a}}{4}\phi_{a}^{4}+\frac{K}{2}|\bm{\nabla}\phi_{a}|^{2}\right]$ (1)
@@
$\chi_{ab}=\chi_{ba}$ is Flory-Huggins type interaction parameter between two species $a$ and $b$. The active part of the chemical potential is
$\mu_{a}^{(\rm ac)}=\sum_{b}\alpha_{ab}\phi_{b},$ (2)
where $\alpha_{ab}=-\alpha_{ba}$.
To make progress, we have made some assumptions while still retaining enough complexity. The interfacial tension is assumed to be identical for all species and equal to $K$, in order to focus on effects that are distinct from a Turing type instability that appears when the coefficients are chosen to be distinct [75]. The coefficient of self-interaction $\Theta_{a}=(1+\Theta)$ is the same for all species, a parameter akin to an external tuning parameter, such as the temperature of the system, that controls the transition to a patterned state. Rescaling time, space, and the scalar fields as $t\to t{K}/{\Gamma}$, $x\to x\sqrt{K}$, and $\phi_{a}\to\phi_{a}/\sqrt{s_{a}}$, the equations can be recast as follows by introducing an interaction matrix $\mathbb{M}$ with elements $M_{ab}=\chi_{ab}+\alpha_{ab}$
@@
$-\nabla^{4}\phi_{a},\,\,\int\mbox{d}^{d}\bm{r}\phi_{a}(\bm{r},t)=\phi_{a0},$ (3)
where the term quartic in space gradients represents the surface tension and appears due to the interfacial contribution to $f$ in Eq. (1). From this point onwards, we consider an ensemble of systems where all elements of $\mathbb{M}$ are drawn from a Gaussian distribution of zero mean and variance $\sigma$, a quantity that sets the scale of the inter-species interactions.
$M_{ab}(\bar{\chi},\bar{\alpha})=\frac{\bar{\alpha}}{2\sqrt{N(\bar{\alpha}^{2}+\bar{\chi}^{2})}}A_{ab}+\frac{\bar{\chi}}{2\sqrt{N(\bar{\alpha}^{2}+\bar{\chi}^{2})}}S_{ab},$ (4)
where $\mathbb{A}$ and $\mathbb{S}$ are fully asymmetric and symmetric matrices respectively, i.e. $A_{ab}=-A_{ba}$, and $S_{ab}=S_{ba}$. Eq. (4) follows from the general rule that any matrix can be expressed as the sum of a symmetric and an antisymmetric matrix. Eq. (4) is written such that the variance of the coefficients is $\sigma^{2}$ and the correlation between elements $M_{ab}$ and $M_{ba}$ of $\mathbb{M}$ determines the level of nonreciprocity parameterized by $\alpha$ which depends on the ratio of $\bar{\alpha}$ and $\bar{\chi}$. The properties of $\mathbb{M}$ are summarized as follows
$\langle M_{ab}M_{ab}\rangle=\frac{\sigma^{2}}{4N},\,\,\langle M_{aa}\rangle=0,$
$\langle M_{ab}M_{ba}\rangle=\frac{\sigma^{2}}{4N}(1-2\alpha^{2}),\,\,\alpha=\frac{\bar{\alpha}\bar{\chi}^{-1}}{\sqrt{\bar{\alpha}^{2}\bar{\chi}^{-2}+1}}.$ (5)
@@
To determine the linear stability of the system we substitute $\phi_{a}=\phi_{a0}+\delta\phi_{a}$ in Eqs. (II) to obtain the following linearised dynamics in Fourier space with wavenumber $q$ retaining terms only quadratic order in $q$
$\partial_{t}\delta\phi_{a}$
$=-q^{2}\sum_{b}\left[M_{ab}+(1+\Theta+3\phi^{2}_{a0})\delta_{ab}\right]\delta\phi_{b}+\mbox{O}(q^{4}).$ (6)
Notice that the parameter $\sigma$ simply renormalizes the effective temperature and can be set to unity without any lack of generalization in all calculations, other than in Section IV where we discuss the threshold of stability using analytical arguments. Defining a diagonal matrix $C_{aa}=3\phi^{2}_{a0}$, whose off-diagonal entries are zero, it is clear from Eq. (II) that the linear stability of the system is determined by the eigenvalues of the matrix $\mathbb{D}=\mathbb{M}+\mathbb{C}$. The effect of varied composition thus enters the dynamical matrix as diagonal disorder through the diagonal matrix $\mathbb{C}$ whose diagonal elements $\bar{\phi}$ are drawn from a distribution $P(\bar{\phi})$, as illustrated in Fig. 1 (a).
Figure 2: Spinodal instability in theory and simulations: (a) The boundary of linear stability is plotted with a solid line in the $\Theta-\alpha$ plane, for a fixed set of parameters for the discrete bimodal distribution, see Eq. (12). The predictions are tested using simulations with 5050 realisations of $\mathbb{M}$. The colour of the markers reflects the ensemble and space-averaged order parameter $\Delta$ [see Eq. (10)], which vanishes outside the spinodal shown by a light hue and changes discontinuously to a finite value inside the unstable region, as marked by a darker hue. The grey line is the same for an identical mean composition of the densities. The plots show that the relative stability depends both on the distribution and on $\alpha$ (b) Same as panel (a) for the exponential distribution.
@@ III Spinodal
The stability of the system has to be determined in a multidimensional space constituted by the the effective temperature $\Theta$ and the parameters defining the distribution $P$ from which the average densities of the species are selected. The boundary of the spinodal instability separates the part of the multi-dimensional parameter space where at least one of the N eigenmode in Eq. (II) is linearly unstable (and grows exponentially) from the region where the homogeneously mixed state is stable. The onset of the spinodal instability is thus determined by the eigenvalue $\lambda_{0}$ of $\mathbb{D}$ with the smallest real part. With this definition, we can discuss the onset of pattern formation in the system as controlled by the effective temperature $\Theta$.
Setting the R.H.S. of Eq. (II) to zero, we determine the the threshold value of $\Theta=\Theta_{\rm sp}$ for the mixed state to be stable to perturbations.
@@
where $\{D\}$ denotes the set of parameters that characterize the distribution $P$. On lowering $\Theta$ below $\Theta_{\rm sp}$, the mixture undergoes a spinodal instability as shown in the kymographs of Fig. 1 (c) to evolve to a patterned final state. The statistics of $\lambda_{0}$ are determined by the properties of the dynamical matrix $\mathbb{D}$. For $\Theta$ lower than $\Theta_{\rm sp}$, the homogeneous state is unstable and patterns are formed in the steady state.
For $P(\bar{\phi})=\delta(\bar{\phi}-\phi_{0})$ , i.e. if the mean density of each species is identical, $\lambda_{0}$ is given by the eigenvalues of $\mathbb{M}$ shifted by a constant factor [39].
@@
Eq. (8) implies that for a fixed $\phi_{0}$, $\Theta_{\rm sp}$ is smaller for the active system in comparison to the passive system ($\alpha=0$). The control parameter $\Theta+1$ is synonymous with the strength of self-interaction. If all entries of the matrix $\mathbb{M}$ were set to zero, the non-interacting system would not phase separate for $\Theta>-1$. $\Theta<-1$ implies that interactions between particles of the similar is attractive and leads to phase separation. Eq. (8) shows that random passive interactions lead to $\Theta_{\rm sp}=0$, meaning that they generate effective attractive interactions between the species, such that to tune the system to a homogeneous state, it is necessary to increase intra-species repulsive interactions. For a fully nonreciprocal system, $\alpha=1$, $\Theta_{\rm sp}$ decreases to 1-1 which is the threshold for the non-interacting system. The nonreciprocal counterpart is thus more stable than the reciprocal system. The factor of unity in Eq. (7) simply sets $\Theta_{\rm sp}$ for passive interactions to zero which serves as a good reference point for comparisons.
@@
The slope $m$ and $l_{0}$ depend only on the parameters D{D} characterizing the distribution $P$; in particular, both of these parameters are independent of $\alpha$. The relation in Eq. (9) is rigorously true in two cases; it is a good estimate for an ensemble of systems with interactions following the statistical rules laid out in Eq. (4) for finite NN, and Eq. (9) is rigorously true for a system with infinitely many components. When $N\to\infty$, the variance of $\mbox{Re}(\lambda_{0})$, which goes down with NN as $1/\sqrt{N}$, approaches zero. For finite NN we expect deviations from the relation in Eq. (9). Keeping this in mind, we choose $N=100$ in most of the simulations that are carried out to verify Eq. (9), as for this value we expect fluctuations of $10\%$ in $\lambda_{0}$ for $\alpha=0$, which converges to 1-1 for $N\to\infty$. The numerical results presented in the paper are also averaged over an ensemble of systems to compensate for the expected variations.
First consider the system with only reciprocal interactions ($\alpha=0$), with any specified $P$. Notice that $l_{0}$ is the only parameter that determines the stability of the passive system in the presence of compositional disorder. For different distributions $P$, we can compute $l_{0}$ and use the results to compare the effect that a particular choice of $P$ has on the stability. The lower the value of $l_{0}$, the more stable the system is, as the transition to a patterned state happens at a lower effective temperature (see discussion after Eq. 8). We now discuss the stability of the uniform mixture on tuning the level of nonreciprocity $\alpha$. The sign of the ‘slope’ $m$ between $\Theta_{\rm sp}$ and $\alpha^{2}$ determines the stability of the active system. For $m<0$ the active system is more stable than the passive one. We will show in Section IV that the slope $m$ is always negative, irrespective of the details of the distribution $P$. This implies a very general consequence of nonreciprocity is the stabilisation of the mixed state. We will also show that the sign of $l_{0}$ depends on the specific choice of $P$.
In Fig. 2 we have plotted the theoretical prediction for $\Theta_{\rm sp}$ for the bimodal and exponential distributions described in IV, see Eqs. (12) and (14). The solid lines in Fig. 2 separate the region where the uniform reference is stable (higher $\Theta_{\rm sp}$) from the unstable region (higher $\Theta_{\rm sp}$). To test the predictions of linear stability analysis, we solve Eq. (II) numerically (see Appendix VIII for details). At the beginning of the time evolution, the scalar fields are uniform with mean $\phi_{a0}$ such that $3\phi_{a0}^{2}$ is drawn from either the bimodal or the exponential distributions. The initial conditions at the beginning of the numerical simulation is thus $\phi_{a}=\phi_{a0}+\delta\phi_{a}$, where small deviations $\delta\phi_{a}$ are uncorrelated in space and drawn from a Gaussian distribution. For other details on system size, methods used, see the Appendix VIII. The results of the simulation are shown in Fig. 2, where the markers denote the points where the simulations are carried out and the colour of the markers carries information about the magnitude of the order parameter.